# Gain-Scheduling Data-Enabled Predictive Control for Nonlinear Systems with Linearized Operating Regions

### v1

@@ 1 Introduction
The same principle can be transferred to DeePC. Instead of constructing a single data representation valid for the entire system range with different nonlinear dynamics, one may construct multiple data sets, each corresponding to a local operating region, and schedule between them online. This leads to a Gain-Scheduled DeePC (GS-DeePC) framework, where region-specific Hankel matrices represent the local linear dynamics, and the controller selects the appropriate data matrices according to the current operating condition. In this way, GS-DeePC inherits the advantages of gain scheduling while remaining fully data-driven and model-free.
Two related approaches address nonlinear DeePC within different frameworks. Verhoek et al. (2021) proposed LPV-DPC, which embeds DeePC into a linear parameter-varying (LPV) formulation using measurable scheduling variables and an LPV extension of the fundamental lemma (Verhoek et al., 2025). This enables nonlinearities to be captured within a single LPV-based data representation. The approach has a strong theoretical foundation but depends on accurate, known scheduling signals, a suitable LPV embedding, and leads to a more complex formulation of DeePC. In the numerical results of this work, LPV-DPC is used as the benchmark.
@@
A third approach, developed independently and published at the same time as our work, also considers region-based extensions of DeePC (Guerrero et al., 2025). While both approaches share the idea of using multiple data sets corresponding to different operating regimes, the present work contributes in several distinct directions. First, we provide a systematic and explicit data-collection methodology, including region partitioning, persistency-of-excitation requirements, and a discussion of the trade-off between region width and data availability. Second, instead of overlap bands and parameter interpolation, we introduce composite Hankel matrices that merge neighboring regions directly at the data level, and dwell-time-based switching rules, both aimed at suppressing chattering and ensuring smooth transitions between linearization regions. Third, we emphasize the computational benefits of regional Hankel matrices and analyze how GS-DeePC reduces data requirements compared to standard DeePC. Finally, we conduct an extensive nonlinear benchmark including a sensitivity analysis over the number of linearization regions and a direct comparison with LPV-DPC.
@@ 2 Notation
$x_{k}\in\mathbb{R}^{n}$ denotes the value of $x$ at the discrete time step $k$, where $\mathbb{R}$ denotes the set of real numbers and $n$ the dimension. $\|x\|_{W}^{2}$ denotes the squared weighted euclidean norm of the vector $x$ with the positive weighting matrix $W$, where $\|x\|_{W}^{2}:=x^{\top}Wx$. A system is partitioned into $M_{n}$ number of regions with $\{M_{i}\}_{i=1}^{M_{n}}=[M_{1},\dots,M_{n}]$ where $M_{i}$ denotes each individual region. A composite region $C_{i}$ represents to neighboring regions $C_{i}=M_{i}\cup M_{i+1}$ where the whole system is partitioned into $C_{n}$ number of regions with $\{C_{i}\}_{i=1}^{C_{n}}=[C_{1},\dots,C_{n}]$.
@@ 3 Data-Enabled Predictive Control
$U_{d}=\begin{bmatrix}U_{p}\\ U_{f}\end{bmatrix},\qquad Y_{d}=\begin{bmatrix}Y_{p}\\ Y_{f}\end{bmatrix}$ (2)
@@
Here, $g\in\mathbb{R}^{T-T_{\mathrm{ini}}-N+1}$ serves as the coefficient vector that reconstructs admissible trajectories from the data, and $r\in\mathbb{R}^{p}$ denotes the reference signal. The weighting matrices $Q\in\mathbb{R}^{p\times p}$ and $R\in\mathbb{R}^{m\times m}$ are positive definite. Regularization terms $\|g\|_{\lambda_{g}}^{2}$ and $\|\sigma_{y}\|_{\lambda_{y}}^{2}$, following (Huang et al., 2023), enhance robustness to noise and modeling imperfections. Input and output limitations are enforced through box constraints and the admissible sets $\mathcal{U}$ and $\mathcal{Y}$. Robustness and stability guarantees for DeePC are established in (Berberich et al., 2020; Berberich and Allgöwer, 2024).
@@ 4 Gain-Scheduled DeePC
The Gain-Scheduled DeePC (GS-DeePC) framework extends the classical DeePC formulation to nonlinear systems by combining multiple locally linear data representations. The operating domain of a scheduling variable is partitioned into several regions, each associated with a Hankel matrix constructed from persistently exciting data in that region.
@@ 4.1 Construction of Regional Hankel Matrices
The data-collection process follows four steps: (i) recording input–output trajectories together with the scheduling variable, (ii) partitioning the operating range and extracting corresponding data regions, (iii) assembling regional Hankel matrices, and (iv) constructing composite regions.
@@ 4.1.1 Step 1: Recording Sequences
Following the standard DeePC setup, we collect an input–output trajectory $\{u(k),y(k)\}_{k=1}^{T}$ and, in addition, record a scheduling variable $\rho(k)$ that parameterizes the dominant nonlinear behavior.
@@ 4.1.2 Step 2: Extracting Regional Data
The observed range of $\rho(k)$ is divided into $M_{n}$ regions
$M_{1},\dots,M_{n},$ (5)
each defined by interval boundaries with the according limits $l\in\mathbb{R}$ in (6), where $l_{1}$ and $l_{M_{n}+1}$ are the lower and upper limits of the overall system, respectively.
$b_{i}:=\{\,\psi\in\mathbb{R}\mid l_{i}\leq\psi\leq l_{i+1}\,\}\;\text{for}\;\;i=1,\dots,M$ (6)
@@
For each region $M_{i}$, we extract all input–output subsequences for which the scheduling variable remains within its boundary $b_{i}$. Only subsequences of length at least $T_{\mathrm{ini}}+N$ are extracted.
@@ 4.1.3 Step 3: Construction of Regional Hankel Matrices
For each region $M_{i}$, Hankel matrices $H_{u,i}$ and $H_{y,i}$ are constructed from all retained subsequences. When multiple subsequences are available, we form a mosaic-Hankel matrix in the sense of (Van Waarde et al., 2020) by concatenating the individual Hankel blocks, resulting in a single comprehensive Hankel matrix per region. Although different regions may yield different sequence lengths, we impose the following for uniformity.
@@ 4.1.4 Step 4: Construction of Composite regions
To reduce sensitivity to boundary transitions, we construct composite regions $C$ by merging neighboring regions as in (9). The data and Hankel matrices of the two constituent regions are concatenated as in (10), resulting in a composite representation that spans the composite boundaries $b_{C,i}$ over both linearization regimes with (11).
$C_{i}=M_{i}\cup M_{i+1}\;\;\text{for}\;\;i=1,\dots,M_{n}-1$ (9)
$\mathscr{H}_{C,i}(u)=\big[\mathscr{H}_{i}(u),\mathscr{H}_{i+1}(u)\big]$ $\text{for}\;i=1,\dots,M-1$ (10)
$\mathscr{H}_{C,i}(y)=\big[\mathscr{H}_{i}(y),\mathscr{H}_{i+1}(y)\big]$ $\text{for}\;i=1,\dots,M-1$
$b_{C,i}=[l_{C,i},l_{C,i+1}]=[l_{i},l_{i+2}]\;\;\;\text{for}\;i=1,\dots,M-1$ (11)
The Hankel matrices of the two neighboring composite regions $C_{i}$ yield the linearized behavior of two regions $M_{i}$ and $M_{i+1}$ in equal parts, sharing one linearized subbehavior with each neighboring composite region $C_{i-1}$ and $C_{i+1}$, resulting in overlapping composite regions.
@@
Feasibility of the GS-DeePC optimization problem is preserved through two mechanisms. First, composite regions provide overlapping local linearizations, increasing the likelihood that the active predictor remains compatible with the initialization during transitions. Second, standard DeePC regularization through $\lambda_{g}$ and $\lambda_{ini}$ ensures feasibility under mild inconsistencies between data sets, weak nonlinearities, or measurement noise (Coulson et al., 2019). Because each (composite) region is constructed from locally consistent data and satisfies Assumption 5, the GS-DeePC controller inherits the practical stability and robustness properties of DeePC within each region.
Remark 2: While a formal global stability proof under switching is beyond the scope of this work, the combination of composite regions and regularization provides a pragmatic mechanism that ensures smooth, stable closed-loop behavior in all numerical experiments.
Assumption 5: The nonlinear system is partitioned into a sufficiently large number of regions $\{M_{i}\}_{i=1}^{M_{n}}$ such that, within each region $M_{i}$ and each composite region $C_{i}$, the input–output data used to construct the corresponding Hankel matrices $H_{C_{i}}$ and $H_{M_{i}}$ is generated by a locally linear or weakly nonlinear dynamics that is well approximated by an LTI representation on that region. Consequently, the DeePC controller operating on $H_{C_{i}}$ or $H_{M_{i}}$ satisfies the theoretical assumptions required in (Coulson et al., 2019; Berberich et al., 2020) (persistence of excitation, data consistency, and feasibility) when constrained to the region boundaries $b_{C,i}$ or $b_{i}$, respectively.
Assumption 5 implicitly requires that each region is chosen sufficiently small such that the underlying nonlinear dynamics can be approximated by a single locally linear behavior. If the regions are too wide, the system may exhibit different nonlinear dynamics within the same region, violating the assumption of local linearity and degrading prediction accuracy. Thus, the choice of $M_{n}$ directly determines the validity of the local linear approximation and is addressed empirically in the numerical section. On the other hand, Assumption 3 requires that each region is sufficiently wide for the system to remain within its boundaries long enough to produce subsequences of length $T_{\mathrm{ini}}+N$. Consequently, Assumptions 3 and 5 impose opposing requirements: narrow regions improve the validity of the local linearization, whereas wider regions facilitate the collection of persistently exciting data. In practice, a balance must be struck between these two effects when selecting the region width $w$ or, equivalently, the number of regions $M_{n}$.
@@ 4.2.3 Composite Region Switching
Remark 4: Guerrero et al. (2025) refer to hard and soft switching, which cover similarities to our region switching and composite-region switching, yet with a different implementation of overlapping regions. Such parallels are expected, as both mechanisms are well established in the gain-scheduling literature (Bett, 2005).
@@ 5 Results
To validate the performance of the proposed GS-DeePC, we introduce the DC-Motor with an unbalanced disc example, representing a strong nonlinear system, and compare it with the established LPV-DPC approach. LPV-DPC was introduced in (Verhoek et al., 2021) in detail, and applied to the exact same nonlinear system. The LPV-DPC used for producing the benchmark is the exact implementation used within (Verhoek et al., 2021). Compared with the proposed GS-DeePC framework, LPV-DPC differs structurally in two key aspects. First, it relies on an LPV embedding of the nonlinear system, requiring auxiliary scheduling trajectories and product terms that increase modeling effort and computational overhead. Second, LPV-DPC incorporates the scheduling variable directly into the Hankel matrices, enabling the linearization to vary at every step of the prediction horizon. In contrast, GS-DeePC employs fixed regional Hankel matrices over the entire horizon, offering a simpler formulation and requiring only local data. Thus, LPV-DPC trades a well-defined and measurable scheduling map for within-horizon variability, while GS-DeePC emphasizes simplicity, locality of data, and practical applicability.
The system under investigation is a simulation model of a DC motor coupled to an unbalanced disc, representing a physical system with inherent nonlinear dynamics. The control objective is reference tracking of the angular position $y(t)=\theta(t)$ and the differential equation is given by (12), where $u$ is the controlling input voltage to the system, and $m,g,l,J,\tau,K_{m}$ are the physical parameters of the system, outlined in table 1 along with the sampling time $T_{s}$.
@@
$\left[\begin{array}[]{l}\dot{x}_{1}\\ \dot{x}_{2}\end{array}\right]=\left[\begin{array}[]{c}-1/\tau\\ 1\end{array}\right]x_{1}+\left[\begin{array}[]{c}-mgl/J\\ 0\end{array}\right]\sin\left(x_{2}\right)+\left[\begin{array}[]{c}K_{m}/\tau\\ 0\end{array}\right]u$ (14)
The input is constrained to $\mathbb{U}:=[-0.25,0.25]$ according to (Verhoek et al., 2021), while the output constraint set was increased to $\mathbb{Y}:=[-\pi,\pi]$ to increase the nonlinear range of the sinus-function compared to (Verhoek et al., 2025).
To obtain enough and sufficiently rich data that explores the full nonlinear operating range, the system is excited by a combination of low-amplitude randomized inputs and a simple boundary-reflection mechanism. The input signal is constructed as the sum of several independently generated random binary sequence (RBS) components, which provide broadband, persistently exciting variations around an evolving state trajectory. Whenever the measured output approaches a predefined upper or lower fraction of the overall range, an additional constant bias term is added to the input, effectively “pushing” the trajectory back toward the other range boundary. This bang–bang-like correction induces repeated upward and downward sweeps across the full range with an overlying persistently excited multi-RBS, ensuring that all operating regions are sufficiently visited and that the resulting input–output data are suitable for constructing the regional and composite Hankel matrices used in GS-DeePC, as introduced in 4.
@@
To evaluate the control performance in a manner aligned with the objectives of GS-DeePC, the tracking error is separated into steady-state and transient components. Specifically, a steady-state RMSE (RMSESS) is computed only over the quasi-stationary intervals $t\in[2,3],[6,7],[10,11],[14,15],[18,19]\mathrm{s}$, whereas a transient RMSE (RMSET) is evaluated outside these intervals. This separation is necessary because the non-smooth step reference induces large transient errors that dominate the overall RMSE and neglect the effect of improved prediction accuracy. Since the primary benefit of GS-DeePC lies in reducing the steady-state tracking error through more accurate local linearizations, splitting the error metrics enables a clearer and more meaningful assessment of the controller’s performance. Also, the relative error $e_{\mathrm{rel,SS}}$ is displayed only for the same quasi-stationary intervals as for (RMSESS) to provide a meaningful graphic evaluation.
A rough hyperparameter (HP) tuning was performed prior to the experiments, revealing a broad performance plateau over a wide range of values for $\lambda_{g}$ and $\lambda_{ini}$. Consequently, the same set of HPs listed in Tab. 2 was used for all regions, as these values consistently lay in the plateau and yielded robust performance across all approaches. A precise, performance-dependent HP tuning for each individual region could improve the results minimally. However, such an automated region-wise tuning procedure lies outside the scope of this work.
First, we analyze the results in Fig. 1 for GS-DeePC using $M_{n,1}=4,M_{n,2}=8$, and $M_{n,3}=16$ composite regions. As the overall operating range is partitioned more finely into a larger number of locally linear regions, the prediction accuracy improves and consequently, the steady-state performance increases. This trend is clearly reflected in the relative steady-state error, where GS-DeePC consistently outperforms standard DeePC. While standard DeePC is in principle capable of controlling the nonlinear system, it does so with a substantially larger steady-state error and requires a much larger Hankel matrix to sufficiently cover the full nonlinear operating range. The HPs used for all methods are summarized in Tab. 2. Compared to LPV-DPC, GS-DeePC with $M_{n,3}=16$ achieves nearly identical performance, with LPV-DPC yielding only marginally smaller steady-state errors. Both approaches, however, operate in a regime of comparably low steady-state error for this experiment. The final subplot in Fig. 1 illustrates the online switching behavior between composite regions. The transitions are robust and stable, confirming that the composite-region construction provides a robust switching mechanism with no observable chattering.
@@
Figure 1: Reference tracking, relative state space error and active composite region for classic DeePC, LPV-DPC, and GS-DeePC with different number of linearization regions $M_{n,1}=4,M_{n,2}=8$, and $M_{n,3}=16$.
@@
Figure 2 reports the steady-state and transient tracking errors as a function of the number of composite regions $C_{n}$. The case $C_{n}=1$ corresponds to standard DeePC, while the rest represent GS-DeePC with increasingly finer partitioning via the total number of composite regions $C_{n}$. Both RMSESS and RMSET improve rapidly when introducing the first few composite regions, after which the performance plateaus with small fluctuations. For $C_{n}>25$, GS-DeePC yields consistently low RMSEs, confirming that multiple local linearizations significantly enhance prediction accuracy. Increasing the number of regions also yields disadvantages as it requires using more of the collected data points of the overall data sequences, since Assumption 3 enforces a minimum subsequence length $N+T_{ini}$ for increasingly shorter regions.
[…]
@@
Figure 2: RMSE for the transient and steady state error intervals in dependence of composite region numbers, where $C_{n}=1$ marks classic DeePC.
@@
Figure 3 investigates GS-DeePC without composite regions and compares three switching strategies: (i) switching with a selective dwell rule that only prevents immediate switching back into the previously active region for $T_{\mathrm{dwell}}=10\cdot T_{s}$ (ii) switching with a fixed dwell time of $T_{\mathrm{dwell}}=10\cdot T_{s}$ blocking after every transition, and (iii) simple region switching without dwell time,. All three controllers use $M_{n}=8$ (non-composite) regions and successfully stabilize the system with acceptable performance, especially when compared to standard DeePC in Fig. 1 tracking the same reference. The reference step values were purposefully placed near region boundaries to induce chattering. Both dwell-time strategies reduce the frequency of region transitions relative to the simple switching approach without dwell time, but neither fully suppresses chattering.
@@
Compared with the composite-region GS-DeePC in Fig. 1, the increased switching activity is clearly disadvantageous. The frequent region changes lead also to visible fluctuations in the system output $y$, as highlighted in the zoomed insets of both Fig. 1 and Fig. 3. These results confirm that composite regions provide substantially smoother transitions and more robust closed-loop behavior than dwell-time-based switching alone.

### v2

@@ 1 Introduction
The same principle can be transferred to DeePC. Instead of constructing a single data representation valid for the entire system range with different nonlinear dynamics, one may construct multiple data sets, each corresponding to a local operating region, and schedule between them online. This leads to a Gain-Scheduled DeePC framework, where region-specific Hankel matrices represent the local linear dynamics, and the controller selects the appropriate data matrices according to the current operating condition. In this way, GS-DeePC inherits the advantages of gain scheduling while remaining fully data-driven and model-free.
Two related approaches address nonlinear DeePC within different frameworks. Verhoek et al. (2021, 2025a) proposed LPV-DPC, which generalizes DeePC for linear parameter-varying (LPV) systems using measurable scheduling variables and an LPV extension of the fundamental lemma (Verhoek et al., 2025b). This enables nonlinearities to be captured within a single LPV-based data representation. The approach has a strong theoretical foundation, but in the context of nonlinear systems, successful application depends on accurate, known scheduling signals and a suitable LPV embedding, which leads to a more complex formulation of DeePC. In the numerical results of this work, LPV-DPC is used as the benchmark.
@@
Furthermore, another approach, developed independently and published at the same time as our work, also considers region-based extensions of DeePC (Guerrero et al., 2025). While both approaches share the idea of using multiple data sets corresponding to different operating regimes, the present work contributes in several distinct directions. First, we provide a systematic and explicit data-collection methodology, including region partitioning, persistency-of-excitation requirements, and a discussion of the trade-off between region width and data availability. Second, instead of overlap bands and parameter interpolation, we introduce composite Hankel matrices that merge neighboring regions directly at the data level, and dwell-time-based switching rules, both aimed at suppressing chattering and ensuring smooth transitions between linearization regions. Third, we emphasize the computational benefits of regional Hankel matrices and analyze how GS-DeePC reduces data requirements compared to standard DeePC. Finally, we conduct an extensive nonlinear benchmark including a sensitivity analysis over the number of linearization regions and a direct comparison with LPV-DPC.
@@ 2 Notation
$x_{k}\in\mathbb{R}^{n_{x}}$ denotes the value of $x$ at the discrete time step $k$, where $\mathbb{R}$ denotes the set of real numbers and $n_{x}$ the dimension. $\|x\|_{W}^{2}$ denotes the squared weighted euclidean norm of the vector $x$ with the positive weighting matrix $W$, where $\|x\|_{W}^{2}:=x^{\top}Wx$. A system is partitioned into $n$ number of regions with $\{M_{i}\}_{i=1}^{n}=[M_{1},\dots,M_{n}]$ where $M_{i}$ denotes each individual region.
@@ 3 Data-Enabled Predictive Control
$\mathscr{H}(u)=\begin{bmatrix}U_{p}\\ U_{f}\end{bmatrix},\qquad\mathscr{H}(y)=\begin{bmatrix}Y_{p}\\ Y_{f}\end{bmatrix}$ (2)
@@
Here, $g\in\mathbb{R}^{T-T_{\mathrm{ini}}-N+1}$ serves as the coefficient vector that reconstructs admissible trajectories from the data, and $r\in\mathbb{R}^{p}$ denotes the reference signal. The weighting matrices $Q\in\mathbb{R}^{p\times p}$ and $R\in\mathbb{R}^{m\times m}$ are positive definite. Regularization terms $\|g\|_{\lambda_{g}}^{2}$ and $\|\sigma_{y}\|_{\lambda_{y}}^{2}$, following (Huang et al., 2023), enhance robustness to noise and modeling inaccuracies. Input and output limitations are enforced through box constraints and the admissible sets $\mathcal{U}$ and $\mathcal{Y}$. Robustness and stability guarantees for DeePC are established in (Berberich et al., 2020; Berberich and Allgöwer, 2024).
@@ 4 Gain-Scheduled DeePC
The GS-DeePC framework extends the classical DeePC formulation to nonlinear systems by combining multiple locally linear data representations. The operating domain of a scheduling variable is partitioned into several regions, each associated with a Hankel matrix constructed from persistently exciting data in that region.
@@ 4.1 Construction of Regional Hankel Matrices
The data-collection process follows four steps: (i) recording input-output trajectories together with the scheduling variable, (ii) partitioning the operating range and extracting corresponding data regions, (iii) assembling regional Hankel matrices, and (iv) constructing composite regions.
@@ 4.1.1 Step 1: Recording Sequences
Following the standard DeePC setup, we collect an input-output trajectory $\{u(k),y(k)\}_{k=1}^{T}$ and, in addition, record a scheduling variable $\rho(k)$ that parameterizes the dominant nonlinear behavior.
@@ 4.1.2 Step 2: Extracting Regional Data
The observed range of $\rho(k)$ is divided into $n$ regions
$\{M_{i}\}_{i=1}^{n}=[M_{1},\dots,M_{n}]$ (5)
each defined by interval boundaries with the according limits $l\in\mathbb{R}$ in (6), where $l_{1}$ and $l_{n+1}$ are the lower and upper limits of the overall system, respectively.
$b_{i}:=\{\,\psi\in\mathbb{R}\mid l_{i}\leq\psi\leq l_{i+1}\,\}\;\text{for}\;\;i=1,\dots,n$ (6)
@@
For each region $M_{i}$, we extract all input-output subsequences for which the scheduling variable remains within its boundary $b_{i}$. Only subsequences of length at least $T_{\mathrm{ini}}+N$ are extracted.
@@ 4.1.3 Step 3: Construction of Regional Hankel Matrices
For each region $M_{i}$, Hankel matrices $\mathscr{H}_{i}(u)$ and $\mathscr{H}_{i}(y)$ are constructed from all retained subsequences. When multiple subsequences are available, we form a mosaic-Hankel matrix in the sense of (Van Waarde et al., 2020) by concatenating the individual Hankel blocks, resulting in a single comprehensive Hankel matrix per region. Although different regions may yield different sequence lengths, we impose the following for uniformity.
@@ 4.1.4 Step 4: Construction of Composite regions
To reduce sensitivity to boundary transitions, we construct composite regions $C$ by merging neighboring regions as in (9). The data and Hankel matrices of the two constituent regions are concatenated as in (10), resulting in $n_{C}=n-1$ composite representations that spans the composite boundaries $b_{C,i}$ over both linearization regimes with (11).
$C_{i}=M_{i}\cup M_{i+1}\;\;\text{for}\;\;i=1,\dots,n_{C}$ (9)
$\mathscr{H}_{C,i}(u)=\big[\mathscr{H}_{i}(u),\mathscr{H}_{i+1}(u)\big]$ $\text{for}\;i=1,\dots,n_{C}$ (10)
$\mathscr{H}_{C,i}(y)=\big[\mathscr{H}_{i}(y),\mathscr{H}_{i+1}(y)\big]$ $\text{for}\;i=1,\dots,n_{C}$
$b_{C,i}=[l_{C,i},l_{C,i+1}]=[l_{i},l_{i+2}]\;\;\;\text{for}\;i=1,\dots,n_{C}$ (11)
The Hankel matrices of the two neighboring composite regions $C_{i}$ yield the linearized behavior of two regions $M_{i}$ and $M_{i+1}$ in equal parts, sharing one linearized local behavior with each neighboring composite region $C_{i-1}$ and $C_{i+1}$, resulting in overlapping composite regions.
@@
Feasibility of the GS-DeePC optimization problem is supported by two mechanisms. First, composite regions provide overlapping local linearizations, increasing the likelihood that the active predictor remains compatible with the initialization during transitions. Second, standard DeePC regularization through $\lambda_{g}$ and $\lambda_{ini}$ ensures feasibility under mild inconsistencies between data sets, weak nonlinearities, or measurement noise (Coulson et al., 2019). Because each (composite) region is constructed from locally consistent data and satisfies Assumption 5, the GS-DeePC controller inherits the practical stability and robustness properties of DeePC within each region.
Remark 2: While a formal global feasibility and stability proof under switching is beyond the scope of this work, the combination of composite regions and regularization provides a pragmatic mechanism that ensures smooth, stable closed-loop behavior in all numerical experiments.
Assumption 5: The nonlinear system is partitioned into a sufficiently large number of regions $n$ such that, within each region $M_{i}$ and each composite region $C_{i}$, the input-output data used to construct the corresponding Hankel matrices $\mathscr{H}_{i}(u)$, $\mathscr{H}_{i}(y)$ and $\mathscr{H}_{C,i}(u)$, $\mathscr{H}_{C,i}(y)$ is generated by a locally linear or weakly nonlinear dynamics that is well approximated by an LTI representation on that region. Consequently, the DeePC controller operating on these local Hankel matrices satisfies the theoretical assumptions required in (Coulson et al., 2019; Berberich et al., 2020) (persistence of excitation, data consistency, and feasibility) when constrained to the region boundaries $b_{i}$ or $b_{C,i}$, respectively.
Assumption 5 implicitly requires that each region is chosen sufficiently small such that the underlying nonlinear dynamics can be approximated by a single locally linear behavior. If the regions are too wide, the system may exhibit different nonlinear dynamics within the same region, violating the assumption of local linearity and degrading prediction accuracy. Thus, the choice of $n$ directly determines the validity of the local linear approximation and is addressed empirically in the numerical section. On the other hand, Assumption 3 requires that each region is sufficiently wide for the system to remain within its boundaries long enough to produce subsequences of length $T_{\mathrm{ini}}+N$. Consequently, Assumptions 3 and 5 impose opposing requirements: narrow regions improve the validity of the local linearization, whereas wider regions facilitate the collection of persistently exciting data. In practice, a balance must be struck between these two effects when selecting the region width $w$ or, equivalently, the number of regions $n$.
@@ 4.2.3 Composite Region Switching
Remark 4: Guerrero et al. (2025) refers to hard and soft switching, which cover similarities to our region switching and composite-region switching, yet with a different implementation of overlapping regions. Such parallels are expected, as both mechanisms are well established in the gain-scheduling literature (Bett, 2005).
@@ 5 Results
To validate the performance of the proposed GS-DeePC, we introduce the DC-Motor with an unbalanced disc example, representing a strong nonlinear system, and compare it with the established LPV-DPC approach. LPV-DPC was introduced in (Verhoek et al., 2021) and further detailed in (Verhoek et al., 2025a) The LPV-DPC used for producing the benchmark is the exact implementation on the same nonlinear system used within (Verhoek et al., 2021). Compared with the proposed GS-DeePC framework, LPV-DPC differs structurally in two key aspects. First, it relies on an LPV embedding of the nonlinear system, requiring auxiliary scheduling trajectories and product terms that increase modeling effort and computational overhead. Second, LPV-DPC incorporates the scheduling variable directly into the Hankel matrices, enabling the linearization to vary at every step of the prediction horizon. In contrast, GS-DeePC employs fixed regional Hankel matrices over the entire horizon, offering a simpler formulation and requiring only local data. Thus, LPV-DPC trades a well-defined and measurable scheduling map for within-horizon variability, while GS-DeePC emphasizes simplicity, locality of data, and practical applicability.
The system under investigation is a simulation model of a DC motor coupled to an unbalanced disc, representing a physical system with inherent nonlinear dynamics. The control objective is reference tracking of the angular position $y(t)=\theta(t)$ and the differential equation is given by (12), where $u$ is the controlling input voltage to the system, and $m,g,l,J,\tau,K_{\mathrm{m}}$ are the physical parameters of the system, outlined in Table 1 along with the sampling time $T_{s}$.
@@
$\left[\begin{array}[]{l}\dot{x}_{1}\\ \dot{x}_{2}\end{array}\right]=\left[\begin{array}[]{c}-1/\tau\\ 1\end{array}\right]x_{1}+\left[\begin{array}[]{c}-mgl/J\\ 0\end{array}\right]\sin\left(x_{2}\right)+\left[\begin{array}[]{c}K_{\mathrm{m}}/\tau\\ 0\end{array}\right]u$ (14)
The input is constrained to $\mathbb{U}:=[-0.25,0.25]$ according to (Verhoek et al., 2021), while the output constraint set was increased to $\mathbb{Y}:=[-\pi,\pi]$ to increase the nonlinear range of the sinus-function compared to (Verhoek et al., 2025b).
To obtain enough and sufficiently rich data that explores the full nonlinear operating range, the system is excited by a combination of low-amplitude randomized inputs and a simple boundary-reflection mechanism. The input signal is constructed as the sum of several independently generated random binary sequence (RBS) components, which provide broadband, persistently exciting variations around an evolving state trajectory. Whenever the measured output approaches a predefined upper or lower fraction of the overall range, an additional constant bias term is added to the input, effectively “pushing” the trajectory back toward the other range boundary. This bang–bang-like correction induces repeated upward and downward sweeps across the full range with an overlying persistently excited multi-RBS, ensuring that all operating regions are sufficiently visited and that the resulting input–output data are suitable for constructing the regional and composite Hankel matrices used in GS-DeePC, as introduced in Section 4.
@@
To evaluate the control performance in a manner aligned with the objectives of GS-DeePC, the tracking error is separated into steady-state and transient components. Specifically, a steady-state RMSE ($RMSE_{\mathrm{SS}}$) is computed only over the quasi-stationary intervals $t\in[2,3],[6,7],[10,11],[14,15],[18,19]\mathrm{s}$, whereas a transient RMSE ($RMSE_{\mathrm{T}}$) is evaluated outside these intervals. This separation is necessary because the non-smooth step reference induces large transient errors that dominate the overall RMSE and neglect the effect of improved prediction accuracy. Since the primary benefit of GS-DeePC lies in reducing the steady-state tracking error through more accurate local linearizations, splitting the error metrics enables a clearer and more meaningful assessment of the controller’s performance. Also, the relative error $e_{\mathrm{rel,SS}}$ is displayed only for the same quasi-stationary intervals as for $RMSE_{\mathrm{SS}}$ to provide a meaningful graphic evaluation.
A rough hyperparameter (HP) tuning was performed prior to the experiments, revealing a broad performance plateau over a wide range of values for $\lambda_{g}$ and $\lambda_{ini}$. Consequently, the same set of HPs listed in Table 2 was used for all regions, as these values consistently lay in the plateau and yielded robust performance across all approaches. A precise, performance-dependent HP tuning for each individual region could improve the results minimally. However, such an automated region-wise tuning procedure lies outside the scope of this work.
First, we analyze the results in Fig. 1 for GS-DeePC using $n_{C,1}=4,n_{C,2}=8$, and $n_{C,3}=16$ composite regions. As the overall operating range is partitioned more finely into a larger number of locally linear regions, the prediction accuracy improves and consequently, the steady-state performance increases. This trend is clearly reflected in the relative steady-state error, where GS-DeePC consistently outperforms standard DeePC. While standard DeePC is in principle capable of controlling the nonlinear system, it does so with a substantially larger steady-state error and requires a much larger Hankel matrix to sufficiently cover the full nonlinear operating range. The HPs used for all methods are summarized in Table 2. Compared to LPV-DPC, GS-DeePC with $n_{C,3}=16$ achieves nearly identical performance, with LPV-DPC yielding only marginally smaller steady-state errors. Both approaches, however, operate in a regime of comparably low steady-state error for this experiment. The final subplot in Fig. 1 illustrates the online switching behavior between composite regions. The transitions are robust and stable, confirming that the composite-region construction provides a robust switching mechanism with no observable chattering.
@@
Figure 1: Reference tracking, relative state space error and active composite region for classic DeePC, LPV-DPC, and GS-DeePC with different number of linearization regions $n_{C,1}=4,n_{C,2}=8$, and $n_{C,3}=16$.
@@
Figure 2 reports the steady-state and transient tracking errors as a function of the number of composite regions $n_{C}$. The case $n_{C}=1$ corresponds to standard DeePC, while the rest represent GS-DeePC with increasingly finer partitioning via the total number of composite regions $n_{C}$. Both $RMSE_{\mathrm{SS}}$ and $RMSE_{\mathrm{T}}$ improve rapidly when introducing the first few composite regions, after which the performance plateaus with small fluctuations. For $n_{C}>25$, GS-DeePC yields consistently low RMSEs, confirming that multiple local linearizations significantly enhance prediction accuracy. Increasing the number of regions also yields disadvantages as it requires using more of the collected data points of the overall data sequences, since Assumption 3 enforces a minimum subsequence length $N+T_{ini}$ for increasingly shorter regions.
For the benchmark LPV-DPC implementation of (Verhoek et al., 2021), the performance obtained in the same experiment was $RMSE_{\mathrm{SS}}=0.002$ and $RMSE_{\mathrm{T}}=0.74$. The best GS-DeePC performance was achieved at $n_{C}=23$ with $RMSE_{\mathrm{SS}}=0.003$ and $RMSE_{\mathrm{T}}=0.74$, showing that GS-DeePC approaches LPV-DPC performance despite its simpler structure. Standard DeePC, in contrast, yields significantly worse results with $RMSE_{\mathrm{SS}}=0.057$ and $RMSE_{\mathrm{T}}=0.76$, illustrating the advantage of using local linearizations over a single global data representation. To provide further intuition for the numerical values, these RMSE metrics in Fig. 2 can be related to the relative steady-state errors shown in Fig. 1, which are based on the same experiment and illustrate the clear qualitative differences.
@@
Figure 2: RMSE for the transient and steady state error intervals in dependence of composite region numbers, where $n_{C}=1$ marks classic DeePC.
@@
Figure 3 investigates GS-DeePC without composite regions and compares three switching strategies: (i) switching with a selective dwell rule that only prevents immediate switching back into the previously active region for $T_{\mathrm{dwell}}=10\cdot T_{s}$, (ii) switching with a fixed dwell time of $T_{\mathrm{dwell}}=10\cdot T_{s}$ blocking after every transition, and (iii) simple region switching without dwell time. All three controllers use $n=8$ (non-composite) regions and successfully stabilize the system with acceptable performance, especially when compared to standard DeePC in Fig. 1 tracking the same reference. The reference step values were purposefully placed near region boundaries to induce chattering. Both dwell-time strategies reduce the frequency of region transitions relative to the simple switching approach without dwell time, but neither fully suppresses chattering.
@@
Compared with the composite-region GS-DeePC in Fig. 1, the increased switching activity is clearly disadvantageous. The frequent region changes also lead to visible fluctuations in the system output $y$, as highlighted in the zoomed insets of both Fig. 1 and Fig. 3. These results confirm that composite regions provide substantially smoother transitions and more robust closed-loop behavior than dwell-time-based switching alone.